In a sample of 300 students, 68% said they own an iPod and a smart phone. Compute a 97% confidence interval for the true percent of students who own an iPod and a smartphone.


1
Expert's answer
2022-01-06T17:35:41-0500

"p'=0.68"

"q'=1-p'=1-0.68=0.32"

Since CL = 0.97

"\alpha=1-0.97=0.03"

"\frac{\alpha}{2}=\frac{0.03}{2}=0.015"

The area to the left of Z is 0.015 and the area to the right of Z is 1 – 0.015 = 0.985.

"Z=2.17"

"CI=p'\pm Z_\frac{\alpha}{2}\sqrt{\frac{p'q'}{n}}"

"CI=0.68\pm 2.17\sqrt{\frac{0.68\times0.32}{300}}"

"CI=0.68\pm 0.0269"

"CI=(0.68-0.0269), (0.68+0.0269)"

"CI=0.6531, 0.7069"

This means 97% confidence interval for the true percent of students who own an iPod and a smartphone is between 0.6531 and 0.7069.